OpenStudy (anonymous):

Find the second derivative if f(x)=x^(1/3)*(x-4)

OpenStudy (anonymous):

probably easier to start with \[f(x)=x^{\frac{4}{3}}-4x^{\frac{1}{3}}\] and use the power rule twice

OpenStudy (anonymous):

I actually missed all of this unit due to migraines so I'd really like to know the power rule!

OpenStudy (anonymous):

So would f''(x)= 4/9(x)^-(2/3) - (8/9)x^(-5/3)? And how do I determine concavity from that?

OpenStudy (amoodarya):

look at this file

OpenStudy (anonymous):

Awesome! Thanks! How do I determine Concavity from that?

OpenStudy (amoodarya):

Set the f''(x) sign now

OpenStudy (anonymous):

Confused by what you mean by that.

OpenStudy (amoodarya):

find the sign of f''(x)

OpenStudy (anonymous):

It would be positive correct?

OpenStudy (amoodarya):



OpenStudy (anonymous):

Ohhhhhhhhhhh. You used the zeros of the second derivative correct?

OpenStudy (amoodarya):